# Trace: Contrastive learning for multi-trial time-series data in neuroscience

### v1

@@ 3.3 Alternative contrastive learning framework for embedding time series
We compare against two other contrastive time-series models, CEED [40] and TS2Vec [43]. CEED provides the closest alternative to our framework, as it uses contrastive learning with general purpose time-series data augmentations and task-specific data augmentations, applicable to extracellular action potential wave forms [40]. To compare with CEED in the context of general neuroscience time series, we re-implemented their general purpose time-series augmentations: (1) amplitude jittering, (2) temporal jittering, and (3) correlated background noise. For more details see Appendix A.6. CEED embeds time series into five-dimensional (5D) space, which is mapped to 2D using PCA for visualization (alternative UMAP results are provided in Fig. S6, Table S3).
TS2Vec employs two types of contrastive losses: In the first, it contrasts representations of different time series (like CEED and Trace). In the second, it contrasts representations of different time points from the same time series. These losses are applied through a hierarchical process – starting with two time segments and progressively subdividing it into smaller units – enabling both inter-series and temporal intra-series contrasting at multiple scales. TS2Vec outputs in $\mathbb{R}^{320}$ and uses the cosine similarity in its loss function, like CEED. When applying CEED and TS2Vec to our neural data, we use the mean neural response across trials. Both methods apply identical transformations for each recorded neuron $i$. In contrast, our approach of using the means across a subset of trials allows to automatically identify the local noise structure for each neuron (Sec. 3.2).
We compared our 2D Trace embeddings with the default versions of CEED and TS2Vec using PCA (as suggested by the CEED paper) to obtain a 2D visualization of their 5D and 320D outputs, respectively. We call these CEED $\rightarrow$ PCA and TS2Vec $\rightarrow$ PCA (Fig. 3d,e). We also ablated Trace’s two main contributions: generating positive pairs by averaging across a subset of trials and Cauchy similarity on 2D outputs. By modifying CEED and TS2Vec to output 2D embeddings with Cauchy similarity, we created hybrid models, Trace $+$ CEED and Trace $+$ TS2Vec (Fig. S3, App. A.5).
@@ 4.1 Datasets
We simulated artificial neuronal responses of two cell types with distinct temporal structure: Each response consisted of 25 seconds of baseline activity followed by 5 seconds of class-specific signal, with either positive or negative amplitude defining the two cell types (see Fig. 2a–d). To mimic realistic recording conditions, we added Gaussian noise with varying intensities: Higher variance during the baseline activity (amplitude 10, standard deviation (SD) from 1 to 60 for single trial) and lower variance during the class-specific response period (amplitude 9 or 11, SD 8 for single trial), resembling known effects of stimulus onset in visual neurons [11]. We generated 10 trials per neuron to mimic typical experimental conditions. While idealized, two neuron types may indeed differ in their response to only part of the stimulus, making long response periods uninformative or even detrimental for separating them.
Large-scale neural dataset
We used a large-scale two-photon-imaging dataset from in vivo mouse retinal ganglion-cell axon endings, measured in superior colliculus of awake, head-fixed mice (Fig. 3a). These neurons expressed the genetic calcium indicator GCaMP8m under the hSyn promoter. We presented the following two visual stimuli to the animal during recordings: (1) a full-field “chirp” consisting of a bright step and two sinusoidal intensity modulations (Fig. 1a), and (2) local bright moving bars on a dark background in eight directions (Fig. 4f, top right). In total, the dataset consisted of recordings of 71,021 individual retinal ganglion cell boutons measured at a sampling frequency of 8 Hz, leading to time series with 260 time bins. For the two stimuli, 15 and 10 repeated trials were recorded, respectively. We manually split responses into four groups using the responses to the moving bar stimulus (see App. A.1, Fig. 3b): ON, OFF, ON-OFF, and Suppressed-by-contrast (Sbc). We used these labels to evaluate the cluster structure of the embeddings. All procedures were approved by the Institutional Animal Care and Use Committee of Anonymous Institution.
@@ 4.2 Quantitative measures of embedding quality
    The ARI score [18] between the response-group labels based on the moving bar time series (ON, OFF, ON-OFF, Sbc, see above) and response groups in the 2D embedding identified using Mixture of Gaussian clustering with 4 clusters. This quantifies how well the 2D embedding captured major response groups.
@@
    The kNN accuracy as a standard metric that quantifies embedding quality by predicting the response-type labels through a majority vote of each point’s 15 nearest neighbors in the 2D embedding space. This metric indicates how well points from the same response group cluster together and is commonly used in contrastive learning [28].
@@
    Spearman’s rank correlation $r_{S}$ [35] between pairwise distances in the original time-series space and the low-dimensional embedding, $r_{S}=\operatorname{corr}(\|x_{i}-x_{j}\|,\|z_{i}-z_{j}\|)$ to quantify how well the embedding preserved the relative distances between data points. This is a common metric for visualization methods [20], but rests on the assumption that distances between the original data are meaningful, which may not be the case due to the curse of dimensionality (Sec. 5.1, [1]).
@@
    The linear or radial correlation $r$ between the 2D embedding and the ON-OFF index ($\text{OOi}\in[-1,1]$), the response-transience index ($\text{RTi}\in[0,1]$) (see App. A.2A.3), and the recording depth in the superior colliculus (Fig. S4). For linear gradients, we first fit a linear regression between the embedding coordinates $(x,y)$ and each variable to determine the rotation angle $\theta$ that maximized correlation along one axis. The radial distance of each point was determined based on its Euclidean distance from the arithmetic mean of the embedding. Absolute Pearson correlations were computed between biological variables and both the $\theta$-direction (linear) and radial distances and the maximum reported in Table 1.
@@ 5 Results
The goal of Trace is to support exploratory data analysis of multi-trial time-series data in neuroscience to identify functional cell types or groups and to study continuous variation of neuronal function between cells. Here, we show that Trace performs better than its competitors both on a synthetic dataset (Fig. 2) and a large-scale dataset of in vivo two-photon-imaging recordings of neural activity in superior colliculus (Fig. 3).
Table 1: Quantitative model performance the large-scale superior colliculus dataset. The columns are: model type (see text), the total number of training epochs, the ARI score, $k$NN accuracy, Spearman correlation ($r_{S}$), the maximum correlation for ON-OFF-index ($r_{\text{OOi}}$), response transience index ($r_{\text{RTi}}$), the recording depth $r_{\text{Depth}}$, and the average rank $\mu_{\text{Rank}}$ of each method. For a definition of the measures, see Sec. 4.2. All metrics other than the rank are better when higher. Uncertainties for the ARI score were insignificant and we omitted them. The best values for each metric are bold.
Model ARI $k$NN acc. $r_{S}$ $r_{\text{OOi}}$ $r_{\text{RTi}}$ $r_{\text{Depth}}$ $\mu_{\text{Rank}}$
@@ 5.1 Trace identifies informative regions in a synthetic dataset
Figure 4: The Trace embedding reflects variables of biological interest and aids in identifying experimental artifacts. (a) Left: Trace embedding color-coded by the ON-OFF index (OOi). Right: Maximum correlation with OOi across all methods. (b) Same as (a), but for the response transience index (RTi), and (c) the recording depth in the superior colliculus. (d) Trace embedding with example islands extracted using HDBSCAN clustering ($\epsilon=0$, minimum numbers of samples per cluster 50, minimum cluster size 500). Each example shows (from left to right) the mean response to the chirp and the moving bar stimulus, the OOi, the RTi, the recorded depth and the anatomical location for the respective cluster. (e) Trace embedding with artifact island (red). (f) Top: Functional characteristics of the artifact island in response to the chirp (left) and the moving bars (right). Bottom: Biological variables (OOi, RTi), depths and anatomical location for responses identified as artifacts. (g) Two example scan fields with neuronal responses identified as artifacts in red and others in gray.
@@ 5.2 Trace represents biological structure better in superior colliculus two-photon dataset
Next, we applied Trace to two-photon calcium recordings of retinal ganglion-cell boutons measured in the mouse superior colliculus responding to light stimuli (Fig. 3a, Sec. 4.1 for details). Based on the recorded time series, we identified four major neural response groups based on the responses to the moving bar stimulus: ON, OFF, ON-OFF, and Suppressed-by-contrast (Sbc) neurons (Fig. 3b). We compared Trace against CEED, TS2Vec, $t$-SNE, and UMAP (Fig. 3c-g, Table 1) and additionally extended the Trace setup to use either CEED-like augmentations or the TS2Vec approach (see Sec. 3.3, Fig. S3). We found that the embedding learned by Trace and Trace-variants visually exhibited the best balance between resolving cluster structure and retaining large-scale structure (Fig. S3. This was confirmed quantitatively: For example, Trace most accurately reflected the manually identified response groups (ON, OFF, ON-OFF, and Suppressed-by-contrast), as indicated by the best ARI score (Table 1). In addition, Trace showed comparable $k$NN accuracy to $t$-SNE, in contrast to CEED or TS2Vec, indicating that nearest neighbour in the embedding typically came from the same neuronal response group. Finally, Trace and Trace-variants showed the best correlation between time-series distances and embedding distances, showing that the embedding overall respected the structure of the high-dimensional space well. Computationally, Trace was significantly faster than TS2Vec (Table S4) and more efficient than CEED. While CEED needs expensive data augmentations for each observation in the batch we pre-computed 10k noise samples to improve efficiency (not counted in the reported training time). This is not necessary for Trace as the mean of a subset of trials is used for the positive pairs. Testing how reducing the number of trials ($k$) in non-overlapping subsets affects the results we found that Trace still achieved good performance even with $k=1$ (Table S2, Fig. S2b).
@@
We next studied to what extent the different embeddings captured other neuronal properties such as the tendency to respond to light increments or decrements (measured by the ON-OFF index, OOi, Fig. 4a) or the kinetics of the response to a light step (measured by the response transience index, RTi, Fig. 4b) [5]. We found that Trace and Trace-variants captured these properties well or better than competing methods, yielding the highest correlation with RTi and comparable correlation values for OOi and depth (Table 1, Fig. 4a-c, Fig. S4). Notably, the depth is a completely independent measure never used during learning (whereas OOi and RTi were derived from the shape of the time-series activity used for training). When ranking methods by their mean rank ($\mu_{\text{Rank}}$) across all evaluation metrics, Trace came first among the evaluated models showing that it consistently performs well across all metrics, in line with the visual impression (Table 1, Fig. 3, Fig. S3).
@@
Next, we clustered the 2D Trace representation using HDBSCAN [25] to explore the structure of the visualization in more detail (Fig. 4d). We found that distinct clusters in the embedding showed unique neural response characteristics, representing types within the manually defined response groups of ON, OFF, ON-OFF, and Sbc (examples shown in Fig. 4d). Interestingly, many of these identified subgroups clustered in specific, spatially distinct regions within the superior colliculus, suggesting that the Trace embedding learned the known relationship between neural responses and anatomical location (without having access to the location during training) [23].
@@ 5.3 Trace finds recording artifacts and outliers
In the Trace embedding, we found an isolated island on the far right, which showed peculiar light responses (Fig. 4e), closely following the light intensity of the stimulus (Fig. 4f). We found that these responses were exclusively recorded on the far left of a scan field (Fig. 4f, g), suggesting that they corresponded to light artifacts. In the representations for Trace $+$ CEED amd CEED this island was also visible but less clearly separated. Interestingly, the only other embedding clearly showing these artifacts as outliers was that of UMAP, while they did not stand out at all for TS2Vec (Fig. S5).

### v2

@@ 3.3 Alternative contrastive learning framework for embedding time series
We compare against two other contrastive time-series models, CEED [vishnubhotla2023towards] and TS2Vec [yue2022ts2vec]. CEED provides the closest alternative to our framework, as it uses contrastive learning with general purpose time-series data augmentations and task-specific data augmentations, applicable to extracellular action potential wave forms [vishnubhotla2023towards]. To compare with CEED in the context of general neuroscience time series, we re-implemented their general purpose time-series augmentations: (1) amplitude jittering, (2) temporal jittering, and (3) correlated background noise. For more details see Appendix A.7. CEED embeds time series into five-dimensional (5D) space, which is mapped to 2D using PCA for visualization (alternative UMAP results are provided in Fig. S6, Table S3).
TS2Vec employs two types of contrastive losses: In the first, it contrasts representations of different time series (like CEED and Trace). In the second, it contrasts representations of different time points from the same time series. These losses are applied through a hierarchical process — starting with two time segments and progressively subdividing it into smaller units — enabling both inter-series and temporal intra-series contrasting at multiple scales. TS2Vec outputs in $\mathbb{R}^{320}$ and uses the cosine similarity in its loss function, like CEED. When applying CEED and TS2Vec to our neural data, we use the mean neural response across trials. Both methods apply identical transformations for each recorded neuron $i$. In contrast, our approach of using the means across a subset of trials allows to automatically identify the local noise structure for each neuron (Sec. 3.2).
We compared our 2D Trace embeddings with the default versions of CEED and TS2Vec using PCA (as suggested by the CEED paper) to obtain a 2D visualization of their 5D and 320D outputs, respectively. We call these CEED $\rightarrow$ PCA and TS2Vec $\rightarrow$ PCA (Fig. 3d,e). We also ablated Trace’s two main contributions: generating positive pairs by averaging across a subset of trials and Cauchy similarity on 2D outputs. By modifying CEED and TS2Vec to output 2D embeddings with Cauchy similarity, we created hybrid models, Trace ++ CEED and Trace ++ TS2Vec (Fig. S3, App. A.6).
@@ 4.1 Datasets
We simulated artificial neuronal responses of two cell types with distinct temporal structure: Each response consisted of 25 seconds of baseline activity followed by 5 seconds of class-specific signal, with either positive or negative amplitude defining the two cell types (Fig. 2a–d). To mimic realistic recording conditions, we added Gaussian noise with higher variance during the baseline activity (amplitude 10, standard deviation (SD) from 1 to 60 for single trial) and lower variance during the class-specific response period (amplitude 9 or 11, SD 8 for single trial), resembling known effects of stimulus onset in visual neurons [churchland2010stimulus]. We generated a typical number of 10 trials per neuron. While idealized, two neuron types may indeed differ in their response to only part of the stimulus, making long response periods uninformative or even detrimental for separating them.
Large-scale neural calcium imaging dataset
We used a large-scale two-photon-imaging dataset from in vivo mouse retinal ganglion cell axon endings, measured in superior colliculus of awake, head-fixed mice (Fig. 3a). These neurons expressed the genetic calcium indicator GCaMP8m under the hSyn promoter. We presented two visual stimuli to the animal during recordings: (1) a full-field “chirp” consisting of a bright step and two sinusoidal intensity modulations (Fig. 1), and (2) local bright moving bars on a dark background in eight directions (Fig. 4f, right). In total, the dataset consisted of recordings of $71,021$ individual retinal ganglion cell boutons measured at a sampling frequency of 8 Hz, leading to time series with 260 time bins. For the two stimuli, 15 and 10 repeated trials were recorded, respectively. We manually split responses into four groups using the responses to the moving bar stimulus (see App. A.1, Fig. 3b): ON, OFF, ON-OFF, and Suppressed-by-contrast (Sbc) and to evaluate the cluster structure of the embeddings we clustered the data using a Gaussian mixture model with 50 components. All procedures were approved by the the Baylor College of Medicine, Houston, USA, animal protocol number: AN-8132.
Allen Institute Neuropixels spiking dataset
We used the Allen Institute Neuropixels “visual coding” dataset [siegle2021survey], which is part of the Allen Brain Observatory to test performance on a Neuropixels spiking dataset with action potential resolution. The dataset recorded activity of single neurons across visual cortical and thalamic structures in awake, head-fixed mice viewing diverse visual stimuli using Neuropixels silicon probes. After quality filtering and excluding non-visual neurons, we analyzed spiking responses of $10,322$ neurons to light flashes and drifting gratings across visual brain areas (for details see App. A.5). To compute the ARI score and kNN accuracy, we used the labels of brain area. Biologically meaningful metrics provided by the Allen Institute were: orientation selective index (OSI), preferred orientation (PO), grating modulation ratio (F1/F0), natural image selectivity (NIS; based on responses to natural images not used to create the embeddings), behavioral modulation (correlation of firing rate with running speed of the mouse; not used to create the embeddings). Alignment of the embedding with biological metrics was measured with kNN regression ($R^{2}$). For PO and behavior, we computed the radial correlation, as it better captured their global structure in the embedding.
@@ 4.2 Quantitative measures of embedding quality
    The ARI score [hubert1985comparing] between response groups identified in the 2D embedding using Mixture of Gaussian clustering and either the group labels of the calcium imaging dataset or brain region of the Neuropixels dataset, respectively. This quantifies how well the 2D embedding captured major response groups.
@@
    The kNN accuracy as a standard metric that quantifies embedding quality by predicting the response-type labels through a majority vote of each point’s 15 nearest neighbors in the 2D embedding space. This metric indicates how well points from the same response group cluster together and is commonly used in contrastive learning [oquab2023dinov2].
@@
    Spearman’s rank correlation $r_{S}$ [spearman1904proof] between pairwise distances in the original time-series space and the low-dimensional embedding, $r_{S}=\operatorname{corr}(\|x_{i}-x_{j}\|,\|z_{i}-z_{j}\|)$ to quantify how well the embedding preserved the relative distances between data points. This is a common metric for visualization methods [kobak2019art], but rests on the assumption that distances between the original data are meaningful, which may not be the case due to the curse of dimensionality ([aggarwal2001surprising]).
@@
    The linear or radial correlation $r$ between the 2D embedding and biological relevant variables such as the ON-OFF index ($\text{OOi}\in[-1,1]$), the response-transience index ($\text{RTi}\in[0,1]$) (see App. A.2A.3), and the recording depth for the calcium imaging dataset (Fig. S4) and preferred orientation (PO) and behavioral modulation for the Neuropixels dataset. For linear gradients, we first fit a linear regression between the embedding coordinates $(x,y)$ and each variable to determine the rotation angle $\theta$ that maximized correlation along one axis. The radial distance of each point was determined based on its Euclidean distance from the arithmetic mean of the embedding. Absolute Pearson correlations were computed between biological variables and both the $\theta$-direction (linear) and radial distances and the maximum reported in Table 1.
@@
    The kNN regression to measure how well the respective biological variable of a neuron is predicted by averaging the values of its $k=15$ nearest neighbors in the embedding (used for the Neuropixels dataset as it sometimes better captured the global structure in the embedding).
  6. 6.
@@ 5 Results
The goal of Trace is to support exploratory data analysis of multi-trial time-series data in neuroscience to identify functional cell types or groups and to study continuous variation of neuronal function between cells. Here, we show that Trace performs better than its competitors both on a synthetic dataset (Fig. 2) and on two large-scale in vivo neuroscience time-series datasets (Figs. 3,S8).
Table 1: Quantitative model performance the large-scale superior colliculus dataset. The columns are: model type (see text), the ARI score, kkNN accuracy, Spearman correlation ($r_{S}$), the maximum correlation for ON-OFF-index ($r_{\text{OOi}}$), response transience index ($r_{\text{RTi}}$), the recording depth $r_{\text{Depth}}$, and the average rank $\mu_{\text{Rank}}$ of each method. For a definition of the measures, see Sec. 4.2. All metrics other than the rank are better when higher. Uncertainties for the ARI score were insignificant and we omitted them. The best values for each metric are bold.
Model ARI kkNN acc. $r_{S}$ $r_{\text{OOi}}$ $r_{\text{RTi}}$ $r_{\text{Depth}}$ $\mu_{\text{Rank}}$
@@ 5.1 Trace identifies informative regions in a synthetic dataset
Figure 4: The Trace embedding reflects variables of biological interest and aids in identifying experimental artifacts. (a) Left: Trace embedding color-coded by the ON-OFF index (OOi). Right: Maximum correlation with OOi across all methods. (b) Same as (a), but for the response transience index (RTi), and (c) the recording depth in the superior colliculus. (d) Trace embedding with example islands extracted using HDBSCAN clustering ($\epsilon=0$, minimum numbers of samples per cluster 50, minimum cluster size 500). Each example shows (from left to right) the mean response to the chirp and the moving bar stimulus, the OOi, the RTi, the recorded depth and the anatomical location for the respective cluster. (e) Trace embedding with artifact island (red). (f) Responses of the artifact island to the chirp (left) and the moving bars (right). (g) Two example scan fields with neuronal responses identified as artifacts in red and others in gray, as well as anatomical location across the entire recording volume.
@@ 5.2 Trace represents biological structure better in superior colliculus two-photon dataset
Next, we applied Trace to two-photon calcium recordings of retinal ganglion cell boutons measured in the mouse superior colliculus responding to light stimuli (Fig. 3a, Sec. 4.1 for details). Based on the recorded time series, we identified four major neural response groups based on the responses to the moving bar stimulus: ON, OFF, ON-OFF, and Suppressed-by-contrast (Sbc) neurons (Fig. 3b). We compared Trace against CEED, TS2Vec, $t$-SNE, and UMAP (Fig. 3c–g, Table 1) and additionally extended the Trace setup to use either CEED-like augmentations or the TS2Vec approach (see Sec. 3.3, Fig. S3). We found that the embedding learned by Trace and Trace-variants visually exhibited the best balance between resolving cluster structure and retaining large-scale structure (Fig. S3). This was confirmed quantitatively: For example, Trace most accurately reflected the manually identified response groups (ON, OFF, ON-OFF, and Suppressed-by-contrast), as indicated by the best ARI score (Table 1). In addition, Trace showed comparable kkNN accuracy to $t$-SNE, in contrast to CEED or TS2Vec, indicating that nearest neighbour in the embedding typically came from the same neuronal response group. Finally, Trace and Trace-variants showed the best correlation between time-series distances and embedding distances, showing that the embedding overall respected the structure of the high-dimensional space well. Computationally, Trace was significantly faster than TS2Vec (Table S5) and more efficient than CEED. While CEED needs expensive data augmentations for each observation in the batch we pre-computed 10k noise samples to improve efficiency (not counted in the reported training time). This is not necessary for Trace as the mean of a subset of trials is used for the positive pairs. Testing how reducing the number of trials ($k$) in non-overlapping subsets affects the results we found that Trace still achieved good performance even with $k=1$ (Table S2, Fig. S2b).
@@
We next studied to what extent the different embeddings captured other neuronal properties such as the tendency to respond to light increments or decrements (measured by the ON-OFF index (OOi), Fig. 4a) or the kinetics of the response to a light step (measured by the response transience index (RTi), Fig. 4b) [baden2016functional]. We found that Trace and Trace-variants captured these properties well or better than competing methods, yielding the highest correlation with RTi and comparable correlation values for OOi and depth (Table 1, Fig. 4a–c, Fig. S4). Notably, the depth is a completely independent measure never used during learning (whereas OOi and RTi were derived from the shape of the time-series activity used for training). When ranking methods by their mean rank ($\mu_{\text{Rank}}$) across all evaluation metrics, Trace came first among the evaluated models showing that it consistently performs well across all metrics, in line with the visual impression (Table 1, Fig. 3, Fig. S3). Trace also performed competitively in higher embedding dimension (Table S4).
@@
Next, we clustered the 2D Trace representation using HDBSCAN [mcinnes2017hdbscan] to explore the structure of the visualization in more detail (Fig. 4d). We found that distinct clusters in the embedding showed unique neural response characteristics, representing types within the manually defined response groups of ON, OFF, ON-OFF, and Sbc (examples shown in Fig. 4d). Interestingly, many of these identified subgroups clustered in specific, spatially distinct regions within the superior colliculus, suggesting that the Trace embedding learned the known relationship between neural responses and anatomical location (without having access to the location during training) [li2023functional].
@@ 5.3 Trace finds recording artifacts and outliers
In the Trace embedding, we found an isolated island on the far right, which showed peculiar light responses (Fig. 4e), closely following the light intensity of the stimulus (Fig. 4f). We found that these responses were exclusively recorded on the far left of a scan field (Fig. 4g), suggesting that they corresponded to light artifacts. In the representations for Trace ++ CEED and CEED this island was also visible but less clearly separated. Interestingly, the only other embedding clearly showing these artifacts as outliers was that of UMAP, while they did not stand out at all for TS2Vec (Fig. S5).
@@
5.4 Trace shows superior performance on Neuropixels spiking dataset
Next, we tested the generalizability of our approach across domains and applied it to the Allen Institute Neuropixels dataset [de2020large, siegle2021survey] that included diverse responses of visual neurons from different brain areas to visual stimuli such as dark and bright flashes and drifting gratings (Fig. S8a).
We evaluated embeddings using distance correlation $r_{S}$ and clustering metrics (ARI score, kkNN accuracy) with brain region as ground truth labels. While clear regional separation is not necessarily expected because hierarchical visual processing naturally creates overlapping response patterns across areas, the brain region provided the only available categorical ground truth for this dataset (Fig. S8b,c). We therefore additionally evaluated embeddings using biologically meaningful continuous variables (OSI, PO, F1/F0; see methods and Table S6). To test generalization beyond the stimulus features used to create embeddings, we also used the natural image selectivity (NIS) and behavioral modulation (correlation with running speed) to evaluate the embedding structure, neither of which were derived from the flash and grating responses used for creating the embeddings (Fig. S9).
While standard methods (t-SNE, UMAP) found some structure especially in terms of distance correlation and for OSI, Trace outperformed them in most metrics and outperformed CEED in all metrics (Table S6; Fig. S8d–h). The performance of TS2Vec variants was mixed, with good preservation of some structure (OSI, NIS) but very poor global distance correlation (Fig. S9). Overall, Trace achieved the best rank by far, indicating best overall performance on this dataset (Table S6). Notably, while Trace was the only methods capturing both image selectivity and modulation by behavior well (despite these features not being used for training), only Trace was able to capture this and produce clearly structured, interpretable embeddings (Fig. S9).